OpenStudy (anonymous):

p=2(L+W) solve for L

OpenStudy (anonymous):

P/2-W=L?

OpenStudy (anonymous):

p=2L+2W Subtract 2W from both sides and you have: 2L = P - 2W : Divide both sides by 2 L = (P-2W)/2 or L = P/2 - W

OpenStudy (anonymous):

lmao.@xoya thanks

OpenStudy (anonymous):

Your welcome @tiffie1234 :)

OpenStudy (anonymous):

T=P+3r/6

OpenStudy (anonymous):

?

OpenStudy (anonymous):

is it 6T-P/3=R @xoya

OpenStudy (anonymous):

T=P+3r/6 solve for R????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

t = (p+3r)/6, multiply by 6 6t = p + 3r, subtract p 6t-p = 3r, divide by 3: (6t-p)/3 = r.

OpenStudy (anonymous):

No doubt you are good at Maths :)

OpenStudy (anonymous):

Hahaha :D okay xD